# Insights from Large-Scale Communications: Initiation

### v1

@@ Abstract
This paper offers an original theoretical framework to quantify the information content associated with cosmological structure formation. The formalism is developed and employed to study the spectrum of information underlying the galaxy distribution in the observable Universe. Using data from SDSS DR18 we further quantify the information sharing across different parts of the studied volume. An attempt to validate the assumption of cosmic homogeneity is made, which rules out the presence of a Universal scale of uniformity below $130{\,h^{-1}\,{\rm Mpc}}$. In addition, an analytical study is carried out to track back the evolution of global information content up to $z=20$, where a log-normal density distribution with redshift-dependent variance, skewness, and kurtosis is used to mimic the observable Universe. A staggering $8\times 10^{144}TB$ of information loss is estimated, caused by the formation of large-scale structures in the present universe. We further illustrate how the global information budget is impacted at different epochs by the interaction between the expansion rate and growth rate of structure, taking the $\Lambda CDM$ model into account. It is found that while the growth rate of the global information content is slowing down, information loss is increasing dramatically despite an ongoing accelerated expansion.
keywords:
large-scale structure of Universe: galaxy clusters-redshift surveys; information theory; methods: analytical-statistical-numerical.
[…]
references.bib
@@ 1 Introduction
The study of galaxy clustering is the cornerstone of modern cosmology. Over the years, it has played an important role in explaining the dynamics of the Universe. peebles73 introduced the primary statistical tools to characterise and quantify the properties of galaxy clustering. Later on, several seminal studies (efstathiou79; hewett82; davis83; bardeen86; kaiser87; blanchard88; hamilton92; landy93; baugh93) refined our understanding of galaxy clustering and its manifestation across different length scales. Both observational facilities and theoretical frameworks have experienced substantial advancements in the last few decades. Large-scale surveys such as the SDSS (york00), 2MASS (skrutskie06), GAMA (driver11) and DES (abbott18) have not only expanded our observational capabilities in terms of individual object detection but also has contributed in understanding the current nature and evolution of galaxy clustering in the Universe. Furthermore, theoretical models and cosmological simulations such as Millennium Run (springel05), Horizon Run (kim15), EAGLE (schaye15) and IllustrisTNG (nelson18) have played a crucial role in interpreting observational data and uncovering the underlying physical mechanisms which govern galaxy clustering in the Universe.
@@
On smaller scales, thermally supported interstellar clouds of self-gravitating gas fragments into stars (jeans1914). In the presence of anisotropic gravitational forces, a much larger, uniformly rotating or non-rotating cold gas cloud undergoes a spheroidal collapse to become a galaxy (lynden62; lin65; binney77). On the other hand, the spherical collapse model (gunn72) explains how large galaxy clusters like Coma can form through the infall of baryonic matter into dark matter (DM) potential wells. However, the formation of the filament and pancake-like structures, that are prevalent in the nearby observable universe, can be explained through the mechanism of anisotropic ellipsoidal collapse (dorosh64; zeldo70; shandarin89). ostriker78 finds evidence of an inside-out dynamical growth of galaxy clustering, which is thought to be originating from the bottom-up hierarchical merging of DM halos (press74; bond91; lacey93; sheth99). Galaxies being the biased tracers of the DM halos are expected to be found in the sharpest density peaks of the DM distribution. To better understand the relationship between a galaxy, its associated DM host, and the space they occupy combined, it is necessary to comprehend the systematic evolution of the information content associated with the configuration of the large-scale structures occupying that space. Long before the galaxy’s formation, this space was characterised by a nearly uniform matter distribution. Over time, gravity forces the matter to collapse and form a bound object, now seen as the galaxy. This process also involves the accretion of additional matter from relatively under-dense surroundings and indirectly helps the under-dense structures to grow. This entire process in every step reduces the randomness of the whole system ( galaxy + surrounding ) and takes it towards a low probability state. The birth of a galaxy requires several conditions to be met; only a few among numerous alternate possibilities would lead to the emergence of the galaxy within a specific space-time span. For example, a larger DM halo nearby could have changed the evolutionary trajectory of the galaxy in terms of space and time. From an alternate point of view, the growth of under-densities due to the combined effect of gravitational drag and the expansion of the Universe creates structures characterised by the enormity of space or nothingness; This again reduces the existing uncertainty of the system. This complex flow of matter and space results in a spatial distribution of probability, and gives rise to a continuum of Information.
@@
Building on the work of nyquist24 and hartley28, shannon48 put in place the groundwork for quantifying the information entropy in classical systems characterised by discrete variables. This led to subsequent research on modelling information exchange within complex systems (jaynes57; wolfram83). hawking76 showed the limitations in information transmission from the ”hidden surface” of a black hole which led beken81 to establish the universal upper bound of entropy. Through dimension reduction, the holographic principle (thooft93; beken94; susskind95) demonstrates that information confined in a finite volume can be represented by information projected through its surface. Furthermore, the relation between the surface area and the upper entropic bound of information holds not just for black holes, but also for vast cosmological scales. Information entropy in inhomogeneous (hosoya04) and homogeneous (pandey19; das19; das23; pandey23) cosmological models has been studied extensively. Several studies validating cosmic homogeneity (pandey15; pandey16; pandey21) and isotropy (sarkar19), using information theory advocate the presence of a Universal scale of uniformity beyond $\sim 200{\,h^{-1}\,{\rm Mpc}}$. Using Fisher information tegmark97 shows how to numerically estimate the accuracy at which observational data can determine cosmological parameters. carron13 presents the sufficient statistics that use the Fisher matrix to measure the information content of a lognormal density distribution originating from a Gaussian initial condition. ensslin09 develop the information field theory for reconstructing spatially distributed large-scale signals. pinho20 quantifies the information content in cosmological probes and connects information entropy to Bayesian inference. Information theory is currently being applied extensively in astronomical data analysis, targeting a lossless recovery of the cosmological information encoded with the incoming signals. Despite this, a comprehensive information-theoretical technique to map the spatial information distribution in the Universe and quantify the amount of information going into creating large-scale structures is still lacking. In this study, we develop a tool that not only allows us to measure the effective amount of spatial information but also helps to determine the information sharing on various length scales associated with overall structure formation. Furthermore, it also aids in presenting a profit and loss statement for the information budget of the Universe.
@@
Throughout the paper, we follow planck18, and use the following set of cosmological parameters for comoving distance calculation, N-body simulation and estimation of global information entropy in $\Lambda CDM$ cosmology: $\Omega_{m}=0.315$;    $\Omega_{\Lambda}=0.685$;    $H_{0}=67.4$;    $\sigma_{8}=0.811$.
@@ 2 Method of Analysis
2.1 Spatial distribution of Information
@@
$\omega_{\mathrm{m}}$ $=$ $\frac{\{(N_{v}-1)+(M-\mathrm{m})-1\}!}{(N_{v}-2)!(M-\mathrm{m})!}.$ (2)
@@
$p\,(\,\mathrm{m}\,)=\frac{\omega_{\mathrm{m}}}{\Omega}\,=\,\frac{\{(N_{v}-1)+(M-\mathrm{m})-1\}!(N_{v}-1)!M!}{(N_{v}+M-1)(N_{v}-2)!(M-\mathrm{m})!},$ (3)
with ${\cal N}=N_{v}+M-1$, Equation 3 becomes
$p\,(\,\mathrm{m}\,)$ $=$ $\left(\frac{N_{v}-1}{{\cal N}-\mathrm{m}}\right)\prod_{i=0}^{\mathrm{m}-1}{\left(\frac{M-i}{{\cal N}-i}\right)}.$ (4)
Figure 1: This figure shows the effect of the approximation proposed in Equation 5
@@
This relation can be used to calculate the probabilities at each voxel in the mesh. However, with $\mathrm{m}$ as an integer, one can only deal with discrete distributions using Equation 4. To evaluate the expected probabilities for any fractional real values of $\mathrm{m}$, one requires an approximation
$\prod_{i=0}^{\mu-1}{\left(\frac{M-i}{{\cal N}-i}\right)}$ $=$ $\left(\frac{M-\frac{\mu}{2}}{{\cal N}-\frac{\mu}{2}}\right)^{\mu};$ (5)
@@
where $\mathrm{m}$ is replaced by $\mu$, a continuous real-valued variable within $0\leq\mu<M$. The objective is to partition any continuous or discrete field into multiple virtual patches, referred to as ”chunks” hereafter, each possessing identical content but varying in morphology. Cumulatively, these chunks cover the entire volume in consideration. After employing the approximation, Equation 4 becomes
$p\,(\,\mu\,)$ $=$ $\left(\frac{N_{v}-1}{{\cal N}-\mu}\right)\left(\frac{M-\frac{\mu}{2}}{{\cal N}-\frac{\mu}{2}}\right)^{\mu}.$ (6)
Now, probability is closely related to information. Probability quantifies the likelihood of an event. On the other hand, information measures the reduction of uncertainty in a system or process through the outcomes. The information content corresponding to the probability of getting $\mu$ objects can be found as,
$I\,(\,\mu\,)$ $=$ $-\log\,\left[\,p(\mu)\,\right]$ (7)
$=$ $\log\left[\left(\frac{{\cal N}-\mu}{N_{v}-1}\right)\left(\frac{2{\cal N}-\mu}{2M-\mu}\right)^{\mu}\right].$
Information is small for events with higher probabilities. On the contrary, unlikely events or configurations hold a large amount of information. The choice of the base of the logarithm here is arbitrary. In this work, we have used natural logarithms; so the information content has unit nats. However, later on the units are converted to terabytes using a conversion factor $u_{I}=\frac{2^{-43}}{\ln 2}$
Let us now introduce the parameter $\gamma$ to control the chunk-to-grid ratio. $\gamma=\frac{M}{N_{v}}$ is effectively the mean number density of chunks. The information content is now obtained as a function of $\mu$ and $\gamma$ by substituting $N_{v}$ with $\gamma$ in Equation 7, i.e.
$I\,(\,\mu\,)=\ln\left[(1+\gamma)-\frac{\gamma(\mu-\gamma)}{M-\gamma}\right]+\mu\ln\left[1+\frac{2(M-\gamma)}{\gamma(2M-\mu)}\right].$ (8)
We can also write $\mu$ in terms of the mean mass fraction $\bar{f}_{\cal M}$, where $\bar{f}_{\cal M}\,=\,\frac{\gamma}{M}$; i.e. the fraction of the total mass contributing to an individual voxel, on average. It also determines the grid resolution. After substituting $M$ with $\frac{\gamma}{\bar{f}_{\cal M}}$ we are left with
$I\,(\,\mu\,)=\ln\left[(1+\gamma)-\frac{(\mu-\gamma)\,\bar{f}_{\cal M}}{(1-\bar{f}_{\cal M})}\right]+\mu\ln\left[1+\frac{1-\bar{f}_{\cal M}}{\left(\gamma-\frac{\mu\,\bar{f}_{\cal M}}{2}\right)}\right].$ (9)
2.2 Information of large-scale Structures
@@
So far in this section, we have not mentioned galaxies or Large-scale Structures. We started with $M$ discrete point objects having unique locations in terms of voxel indices. However, using further approximations, we now have $M$ extended chunks that span throughout the volume in consideration. Hence, they can partially contribute to more than one voxel. Now, let us consider $G$ galaxies distributed in the cubic volume, comprised of $N_{v}$ grids. We can virtually divide $G$ galaxies into $M$ chunks of equal weightage, each having $Q_{M}=\frac{G}{M}$ galaxies. If any given voxel contains $g$ galaxies, we will have $\mu=\frac{g}{Q_{M}}$.
Note that $(\frac{\mu}{\gamma}-1)$ is the fractional density contrast ($\delta$), at any given voxel, and equation Equation 9 can also be written in terms of $\delta$ as
$I(\delta)=\ln\left[(1+\gamma)-\frac{\gamma\delta\bar{f}_{\cal M}}{1-\bar{f}_{\cal M}}\right]+\gamma(1+\delta)\ln\left[1+\frac{(1-\bar{f}_{\cal M})}{\gamma\left\{1-(1+\delta)\frac{\bar{f}_{\cal M}}{2}\right\}}\right].$ (10)
@@
$I(\delta)$ is the information associated with overdensity $\delta$ in the given matter distribution. Equation 10 displays a crucial relationship between the density perturbation and the Information associated with the large-scale structures. Hereafter any mention of information would mean this large-scale information. Equation 10 applies to both discrete and continuous fields since it no longer contains any explicit constraint of discreteness. When we started with $M$ and $N_{v}$ as integers, we could work only with discrete point data. However, one can now think of a continuous field that spans the entire volume but is virtually discretized into $M$ chunks of equal contribution. Equation 10 allows to work with extensive variables that contribute fractionally to the elementary volume. Nonetheless, one has to carefully choose ${\bar{f}_{\cal M}}$ and $\gamma$ depending on the clustering strength ($\delta$) of the distribution. The two terms in Equation 10 on which the logarithm is being operated must produce non-zero positive numbers. This can be easily taken care of by choosing sufficiently large $N_{s}$ or very small values of ${\bar{f}_{\cal M}}$. We also have to remember that the approximation presented in Equation 5 allows us to have the final form of Equation 10, and one has to be careful while using such an approximation. The limitation of this approximation is shown in Figure 1. For $\gamma<0.01$ and $M>100$ one can safely use the approximation with more than $99.9\%$ accuracy up to $\delta\leq 10^{4}$. However, for larger values of $\gamma$, it is necessary to have a large number of chunks in use.
@@
Figure 2: This figure shows probability [left], information [middle] and information gain [right] as a function of $\delta$ for different values of $\gamma$ and $\bar{f}_{\cal M}$. The two panels on the right show the identity inside the square brackets in Equation 15. Symmetric logarithmic scales are used along the vertical axis where the region with the purple shade has a linear scaling.
For a completely homogenous distribution, we would have $\delta=0$ at every grid. If $\bar{I}$ is the average information for a homogeneous distribution, then
$\bar{I}=\ln\left(1+\gamma\right)+\gamma\ln\left[1+\frac{1-\bar{f}_{\cal M}}{\gamma\left(1-\frac{\bar{f}_{\cal M}}{2}\right)}\right]$ (11)
is associated with the base probability of $\bar{p}=e^{-\bar{I}}$.

### v2

# Impact of Large-Scale Structure Formation on the Global Information Budget

@@ Abstract
This paper offers an original theoretical framework to quantify the information content associated with cosmological structure formation. By taking into account the growth in perturbations for the matter density field, we are able to calculate the relative reduction in the information entropy from the projection for a fully homogeneous setting. An analytical study is carried out to track back the evolution of global information content up to $z=20$, where a log-normal density distribution with redshift-dependent variance, skewness, and kurtosis is used to mimic the observable Universe. We report a $0.004\,\%$ drop in the information entropy of the Universe, that is caused by the formation of large-scale structures in the present universe. The accelerated expansion of the Universe is suspected as a countermeasure by nature to override the exponential growth in this entropic drop. We also analyse how the scale factor, expansion rate and growth rate of structure formation are connected with the entropic drop, taking the $\Lambda CDM$ model into account. The formalism is further developed and employed to study the spectrum of information underlying the galaxy distribution in the observable Universe. Using data from SDSS DR18 we also quantify the information sharing across different parts of the studied volume. An attempt to validate the assumption of cosmic homogeneity is made, which rules out the presence of a Universal scale of uniformity below $130{\,h^{-1}\,{\rm Mpc}}$.
@@ 1 Introduction
The study of galaxy clustering is the cornerstone of modern cosmology. Over the years, it has played an important role in explaining the dynamics of the Universe. Peebles(1973) [1] introduced the primary statistical tools to characterise and quantify the properties of galaxy clustering. Later on, several seminal studies [2, 3, 4, 5, 6, 7, 8, 9, 10] refined our understanding of galaxy clustering and its manifestation across different length scales. Both observational facilities and theoretical frameworks have experienced substantial advancements in the last few decades. Large-scale surveys such as the SDSS [11], 2MASS [12], GAMA [13] and DES [14] have not only expanded our observational capabilities in terms of individual object detection but also has contributed in understanding the current nature and evolution of galaxy clustering in the Universe. Furthermore, theoretical models and cosmological simulations such as Millennium Run [15], Horizon Run [16], EAGLE [17] and IllustrisTNG [18] have played a crucial role in interpreting observational data and uncovering the underlying physical mechanisms which govern galaxy clustering in the Universe.
@@
On smaller scales, thermally supported interstellar clouds of self-gravitating gas fragments into stars [19]. In the presence of anisotropic gravitational forces, a much larger, uniformly rotating or non-rotating cold gas cloud undergoes a spheroidal collapse to become a galaxy [20, 21, 22]. On the other hand, the spherical collapse model [23] explains how large galaxy clusters like Coma can form through the infall of baryonic matter into dark matter (DM) potential wells. However, the formation of the filament and pancake-like structures, that are prevalent in the nearby observable universe, can be explained through the mechanism of anisotropic ellipsoidal collapse [24, 25, 26]. Ostriker(1978) [27] finds evidence of an inside-out dynamical growth of galaxy clustering, which is thought to be originating from the bottom-up hierarchical merging of DM halos [28, 29, 30, 31]. Galaxies being the biased tracers of the DM halos are expected to be found in the sharpest density peaks of the DM distribution. To better understand the relationship between a galaxy, its associated DM host, and the space they occupy combined, it is necessary to comprehend the systematic evolution of the information content associated with the configuration of the large-scale structures occupying that space. Long before the galaxy’s formation, this space was characterised by a nearly uniform matter distribution. Over time, gravity forces the matter to collapse and form a bound object, now seen as the galaxy. This process also involves the accretion of additional matter from relatively under-dense surroundings and indirectly helps the under-dense structures to grow. This entire process in every step reduces the randomness of the whole system ( galaxy + surrounding ) and takes it towards a low probability state. The birth of a galaxy requires several conditions to be met; only a few among numerous alternate possibilities would lead to the emergence of the galaxy within a specific space-time span. For example, a larger DM halo nearby could have changed the evolutionary trajectory of the galaxy in terms of space and time. The cosmic-web [32] connects the galaxy clusters through filaments of galaxies that lie on the intersection of walls encompassing voids of various sizes [33, 34, 35]. Hence, any movement of matter across space would cause perturbations in both the spatial distribution of probability and the continuum of information.
@@
Building on the work of Nyquist(1924) [36] and Hartley(1928) [37], Shannon(1948) [38] put in place the groundwork for quantifying the information entropy in classical systems characterised by discrete variables. This led to subsequent research on modelling information exchange within complex systems [39, 40]. Information theory has long been used in astronomical data analysis. Using Fisher information Tegmark(1997) [41, 42] showed how to numerically estimate the accuracy at which observational data can determine cosmological parameters. Later on, Neyrinck(2009) [43] compared the information contents in the power spectra of the density and log-density fields. Concurrently, the information field theory [44] was presented, reconstructing spatially distributed large-scale signals. Pandey(2013) [45] used Shannon entropy to develop the tool for testing the cosmological principle which was utilized in several studies to validate the cosmic homogeneity [46, 47] and isotropy [48]. The Sufficient statistics to quantify the information content of a log-normal density distribution resulting from a Gaussian initial condition was developed in a contemporary work [49]. Recently Pinho (2020) [50] quantified the information content in cosmological probes connecting information entropy to Bayesian inference. Information theory is currently being applied extensively in astronomical data analysis, targeting a lossless recovery of the cosmological information encoded with the incoming signals. Despite these efforts, a comprehensive information-theoretical technique to map the spatial information distribution of the Universe is still not in place. We develop a distinctive identity, the entropic gain that quantifies the fraction of existing information in the Universe that goes into creating large-scale structures. Furthermore, the results presented in the paper aid in understanding the systematic change in the information content of the Universe throughout different redshifts. The evolution of Information entropy in inhomogeneous [51] and homogeneous [52, 53, 54, 55] cosmological models has been extensively studied so far. In this work, however, we provide an entirely new methodology to analyze the entropic change of information resulting from the growth of perturbations in the matter density field, starting from a basic premise. In this study, we develop tools that not only allow us to measure the effective amount of spatial information but also help to determine the information sharing on various length scales associated with overall structure formation. This allows us to investigate the existence of a certain scale of homogeneity in the observable galaxy distribution. The cosmological principle (CP) that advocates the uniformity of the Universe on large scales, is a foundational assumption of modern cosmology that acts as a basis for the standard cosmological model. Several studies in the past have reported that the galaxy distribution in the observable Universe exhibits a fractal nature on small scales but gradually tends towards homogeneity beyond $50-90{\,h^{-1}\,{\rm Mpc}}$[56, 57, 58, 59, 60, 61, 62]. However, several studies suggest this fractal hierarchy [63, 64, 65, 66, 67] or shell-like distribution of galaxy groups[68] persist beyond $120{\,h^{-1}\,{\rm Mpc}}$. Clowes et al (2013) [69] reports the existence of a Large Quasar Group (LQG) extending $>500{\,h^{-1}\,{\rm Mpc}}$; this questions the possibility of the cosmic homogeneity to lie anywhere around a hundred megaparsecs. However, subsequent studies [70, 71, 72] overrule this conclusion by reporting the scale of homogeneity in Luminous Red Galaxy (LRG) and quasar distributions to lie around $150-250{\,h^{-1}\,{\rm Mpc}}$. Despite this, one cannot deny the presence of massive cosmic structures like the Sloan Great Wall[73], Eridanus void[74] or the Local hole[75], in the observable Universe; this again raises questions on the assumption of cosmic homogeneity. As the opinions are still very much divided on the topic of cosmic uniformity, we try to shed light on this issue by using a novel estimator, namely the mutual gain.
@@
Throughout the paper, we adopt the $\Lambda CDM$ cosmology and use the following set of cosmological parameters [76] for comoving distance calculation, N-body simulation and estimation of global information entropy: $\Omega_{m0}=0.315$;    $\Omega_{\Lambda}=0.685$;    $H_{0}=67.4$;    $\sigma_{8}=0.811$.
@@ 2 Method of Analysis
2.1 Probability associated with perturbations in matter density field
@@
$\omega_{\mathrm{m}}$ $=$ $\frac{\{(N_{v}-1)+(M-\mathrm{m})-1\}!}{(N_{v}-2)!(M-\mathrm{m})!}.$ (2.2)
@@
$p\,(\,\mathrm{m}\,)=\frac{\omega_{\mathrm{m}}}{\Omega}\,=\,\frac{\{(N_{v}-1)+(M-\mathrm{m})-1\}!(N_{v}-1)!M!}{(N_{v}+M-1)(N_{v}-2)!(M-\mathrm{m})!},$ (2.3)
with ${\cal N}=N_{v}+M-1$, Equation 2.3 becomes
$p\,(\,\mathrm{m}\,)$ $=$ $\left(\frac{N_{v}-1}{{\cal N}-\mathrm{m}}\right)\prod_{i=0}^{\mathrm{m}-1}{\left(\frac{M-i}{{\cal N}-i}\right)}.$ (2.4)
Figure 1: This figure shows the effect of the approximation proposed in Equation 2.5
@@
This relation can be used to calculate the probabilities at each voxel in the mesh. However, with $\mathrm{m}$ as an integer, one can only deal with discrete distributions using Equation 2.4. To evaluate the expected probabilities for any fractional real values of $\mathrm{m}$, one requires an approximation
$\prod_{i=0}^{\mu-1}{\left(\frac{M-i}{{\cal N}-i}\right)}$ $=$ $\left(\frac{M-\frac{\mu}{2}}{{\cal N}-\frac{\mu}{2}}\right)^{\mu};$ (2.5)
@@
where $\mathrm{m}$ is replaced by $\mu$, a continuous real-valued variable within $0\leq\mu<M$. The objective is to partition any continuous or discrete field into multiple virtual patches, referred to as “chunks” hereafter, each possessing identical content but varying in morphology. Cumulatively, these chunks cover the entire volume in consideration. After employing the approximation, Equation 2.4 becomes
$p\,(\,\mu\,)$ $=$ $\left(\frac{N_{v}-1}{{\cal N}-\mu}\right)\left(\frac{M-\frac{\mu}{2}}{{\cal N}-\frac{\mu}{2}}\right)^{\mu}.$ (2.6)
Let us now introduce the parameters $\gamma$ and $\bar{f}_{\cal M}$ such that $\gamma=\frac{M}{N_{v}}$ is the mean chunk density and $\bar{f}_{\cal M}=\,\langle\frac{\mu}{M}\rangle=\,\frac{1}{Ng}$ is the mean chunk fraction. Also, $\mu$ can be written in terms of fractional density contrast $\delta$ as $\mu=(1+\delta)\gamma$. As there are ${\bar{f}}_{\cal M}^{-1}$ voxels in the entire cube, without prior knowledge of the density profile the probability of finding the density contrast $\delta$ for a given voxel is
[…]
2.2 Information associated with density perturbations
@@
So far in this section, we have not mentioned galaxies or Large-scale Structures. We started with $M$ discrete point objects having unique locations in terms of voxel indices. However, using further approximations, we now have $M$ extended chunks that span throughout the volume in consideration. Hence, they can partially contribute to more than one voxel. Now, let us consider $G$ galaxies distributed in the cubic volume, comprised of $N_{v}$ grids. We can virtually divide $G$ galaxies into $M$ chunks of equal weightage, each having $\frac{G}{M}$ galaxies; Hence, for any given voxel containing $g$ galaxies, the density contrast can be found as $\delta=\frac{g}{G\,\bar{f}_{\cal M}}-1$.
Equation 2.7 exhibits the probability of finding the overdensity $\delta$. Now, probability is closely related to information. Probability quantifies the likelihood of an event and information measures the reduction of uncertainty in a system or process through the outcomes. The information content corresponding to the probability of getting overdensity $\delta$ can be found as,
$I\left(\,\delta\,\right)=\ln\left[(1+\gamma)-\frac{\gamma\delta\bar{f}_{\cal M}}{1-\bar{f}_{\cal M}}\right]+\gamma(1+\delta)\ln\left[1+\frac{(1-\bar{f}_{\cal M})}{\gamma\left\{1-(1+\delta)\frac{\bar{f}_{\cal M}}{2}\right\}}\right].$ (2.8)
@@
Here $I(\,\delta\,)$ is measured in the units of natural bits or nats. Equation 2.8 displays a crucial relationship between the density perturbation and the Information associated with the large-scale structures. Hereafter any mention of information would mean this large-scale information. Equation 2.8 applies to both discrete and continuous fields since it no longer contains any explicit constraint of discreteness. When we started with $M$ and $N_{v}$ as integers, we could work only with discrete point data. Now, one can think of a continuous field that spans the entire volume but is virtually discretized into $M$ chunks of equal contribution. Equation 2.8 allows to work with extensive variables that contribute fractionally to the elementary volume. Also, the shape of the entire region and the elementary volumes are no longer restricted to a specific geometry. However, one has to carefully choose ${\bar{f}_{\cal M}}$ and $\gamma$ depending on the clustering strength ($\delta$). The two terms in Equation 2.8 on which the logarithm is being operated must produce non-zero positive numbers. This can be easily taken care of by choosing sufficiently large $N_{s}$ or very small values of ${\bar{f}_{\cal M}}$. We also have to remember that the approximation presented in Equation 2.5 allows us to have the final form of Equation 2.8, and one has to be careful while using such an approximation. The limitation of this approximation is shown in Figure 1. For $\gamma<0.01$ and $M>100$ one can safely use the approximation with more than $99.9\%$ accuracy up to $\delta\leq 10^{4}$. However, for larger values of $\gamma$, it is necessary to have a large number of chunks in use.
@@
Figure 2: This figure shows probability [left], information [middle] and entropic gain [right] as a function of $\delta$ for different values of $\gamma$ and $\bar{f}_{\cal M}$. The two panels on the right show the identity inside the square brackets in Equation 2.13. Symmetric logarithmic scales are used along the vertical axis where the region with the purple shade has a linear scaling.
For a completely homogenous distribution, we would have $\delta=0$; also for any voxel with a density same as the critical density of the Universe would have the information
[…]
where the associated probability is $\bar{p}=e^{-\bar{I}}$. The information offset from the homogeneous background can now be found as
$\Delta I(\,\delta\,)$ $=$ $I(\,\delta\,)\,-\,\bar{I}$ (2.10)
The increment in information ($\Delta I$) at any given point originates from the relative change in the mass distribution of that point and the surrounding. It is reflected through the reduction in the probability ($\Delta p$) as $\Delta I=-\ln\left[1+\frac{\Delta p}{\bar{p}}\right]$. Hence, a -ve $\Delta p$ would result in a +ve $\Delta I$ and vice-versa.